PYQSidePowered byPrepblink ExamsMy practice

Probability (Class 12)

JEE Main 2025 (Online) 22nd January Evening ShiftMCQMedium📈 Topic in 8% papers+4−1
Ques. 1JEE Main 2025, 22 January (Evening Shift)

If A and B are two events such that P(A B)=0.1, and P(A B) and P(B A) are the roots of the equation 12x2-7x+1=0, then the value of P( A B)P( A B) is :

  1. 43
  2. 74
  3. 94
  4. 53

Answer

(3) 94

Solution

12x2-7x+1=0 x= 13, 14; so P(A),P(B)= 0.11/3, 0.11/4=0.3, 0.4

[figure]

P(A B)=0.3+0.4-0.1=0.6

P( A B)=1-P(A B)=0.9

P( A B)=1-P(A B)=0.4

ratio = 94

Answer: (3)

Keep practising — Probability (Class 12)

Ques. 2JEE Main 2026, 8 April (Evening Shift)

A candidate has to go to the examination centre to appear in an examination. The candidate uses only one means of transportation for the entire distance out of bus, scooter and car. The probabilities of the candidate going by bus, scooter and car, respectively, are 25, 15 and 25. The probabilities that the candidate reaches late at the examination centre are 15, 13 and 14 if the candidate uses bus, scooter and car, respectively. Given that the candidate reached late at the examination centre, the probability that the candidate travelled by bus is :

  1. 1137
  2. 1237
  3. 1337
  4. 1437
Ques. 3JEE Main 2026, 6 April (Morning Shift)

A bag contains (N+1) coins - N fair coins, and one coin with 'Head' on both sides. A coin is selected at random and tossed. If the probability of getting 'Head' is 916, then N is equal to :

  1. 5
  2. 7
  3. 8
  4. 9
Ques. 4JEE Main 2026, 5 April (Morning Shift)

A letter is known to have arrived by post either from KANPUR or from ANANTPUR. On the envelope just two consecutive letters AN are visible. The probability, that the letter came from ANANTPUR, is:

  1. 710
  2. 1017
  3. 1219
  4. 719
Ques. 5JEE Main 2026, 5 April (Evening Shift)

The probabilities that players A and B of a team are selected for the captaincy for a tournament are 0.6 and 0.4, respectively. If A is selected the captain, the probability that the team wins the tournament is 0.8 and if B is selected the captain, the probability that the team wins the tournament is 0.7. Then the probability, that the team wins the tournament, is :

  1. 0.74
  2. 0.76
  3. 0.72
  4. 0.78
Ques. 6JEE Main 2026, 28 January (Morning Shift)

A bag contains 10 balls out of which k are red and (10 − k) are black, where 0 k 10. If three balls are drawn at random without replacement and all of them are found to be black, then the probability that the bag contains 1 red and 9 black balls is:

  1. 7110
  2. 711
  3. 755
  4. 1455
Ques. 7JEE Main 2026, 28 January (Evening Shift)

The probability distribution of a random variable X is given below : array|c|c|c|c|c|c|c|c|c| X & 4k & 307k & 327k & 347k & 367k & 387k & 407k & 6k \ P(X) & 215 & 115 & 215 & 15 & 115 & 215 & 15 & 115 \ array If E(X) = 26315, then P(X < 20) is equal to :

  1. 35
  2. 1415
  3. 815
  4. 1115
Ques. 8JEE Main 2026, 24 January (Morning Shift)

From a lot containing 10 defective and 90 non-defective bulbs, 8 bulbs are selected one by one with replacement. Then the probability of getting at least 7 defective bulbs is

  1. 73108
  2. 67108
  3. 7107
  4. 81108
Ques. 9JEE Main 2026, 23 January (Evening Shift)

Bag A contains 9 white and 8 black balls, while bag B contains 6 white and 4 black balls. One ball is randomly picked up from the bag B and mixed up with the balls in the bag A. Then a ball is randomly drawn from the bag A. If the probability, that the ball drawn is white, is pq, (p, q) = 1, then p + q is equal to

  1. 24
  2. 22
  3. 23
  4. 21
Ques. 10JEE Main 2026, 22 January (Morning Shift)

If a random variable x has the probability distribution arrayccccccccc x & 0 & 1 & 2 & 3 & 4 & 5 & 6 & 7 \ P(x) & 0 & 2k & k & 3k & 2k2 & 2k & k2 + k & 7k2 array then P(3 < x 6) is equal to

  1. 0.34
  2. 0.64
  3. 0.22
  4. 0.33
Ques. 11JEE Main 2026, 21 January (Evening Shift)

A random variable X takes values 0, 1, 2, 3 with probabilities 2a+130, 8a-130, 4a+130, b respectively, where a, b R. Let and respectively be the mean and standard deviation of X such that 2 + 2 = 2. Then ab is equal to:

  1. 12
  2. 60
  3. 30
  4. 3

All Probability (Class 12) PYQs →

PDFDownload Probability (Class 12) PYQs as PDF Free, with answer key and a QR code to every solution — print it or study offline⬇
All JEE Main Mathematics chapters